# Multi-Agent Reinforcement Learning for Connected and Automated Vehicles Control: Recent Advancements and Future Prospects

### v1

@@ III-B Two-dimensional cooperation
Compared to one-dimensional cooperation, two-dimensional cooperation addresses problems related to both longitudinal and lateral control, with its most typical application being cooperative lane changing. As shown in Fig. 7, CAVs communicate and coordinate through V2V and V2I communication channels, enabling them to make decisions and acquire effective lane-changing strategies within shared driving environments.
@@
Hou et al. devise a decentralized cooperative lane-changing controller employing a Multi-Agent Proximal Policy Optimization (MAPPO) approach, which empowers each vehicle to independently learn and assess its policy and action rewards based on local information, while still having access to global state data. The trained policies can be effectively transferred and applied across various traffic conditions, enhancing traffic throughput from uncongested to highly congested scenarios [100]. In [103], the authors put forth a twin-vehicle cooperative driving approach based on the QMIX algorithm. Their method harnesses RL for real-time adaptation in dynamic highway scenarios, striking a balance between autonomy and cooperation via a fair cooperation algorithm, which enables independent lane changing and overtaking in heavy traffic conditions and ensures a fixed formation. These capabilities enable the twin vehicles to boost their speed while skillfully avoiding rigid traffic interference. An enhanced QMIX algorithm [114] is further proposed to enhance the flexibility and effectiveness of the collaborative lane-changing system. They incorporate a stable estimation technique to mitigate the influence of overestimated joint Q-values among agents, which finely balance between maintaining formation and permitting unhindered overtaking, thereby enabling intelligent adaptation to diverse scenarios including heavy traffic, loose traffic, and emergency situations. In an effort to enhance mobility within complex traffic environments, in [115], instead of relying on relative distance and semantic maps, the authors suggest the utilization of traffic states that encapsulate the spatio-temporal interactions between neighboring vehicles. Within the MADRL framework, three prediction models, namely the transformer-based (TS-Transformer), generative adversarial network-based (TS-GAN), and conditional variational autoencoder-based (TS-CVAE) models, are developed and compared for traffic state prediction.
@@
Graph neural networks (GNNs) [116] have gained significant attention in MARL settings, particularly for cooperative lane-changing applications. In the study by Chen et al. [117], they introduce a novel algorithm based on MADQN, which combines a GCN with a deep Q-network. This approach facilitates effective information fusion and decision-making within the MARL framework, thereby enhancing both safety and mobility in cooperative lane-changing scenarios. Moreover, in [118], the authors leverage GNN in conjunction with MADDPG in their work for multi-agent training, addressing the complexity of real-world inputs and aiding in congestion mitigation efforts.
@@
Safety is of paramount importance in the context of CAV scenarios. In [119], the authors have introduced a safety-enhancing actor-critic algorithm that incorporates two innovative techniques. The first technique is the ‘truncated Q-function’, which effectively leverages shared information from neighboring CAVs, ensuring scalability for large-scale CAV systems. The second technique, “safe action mapping”, provides safety guarantees throughout both the training and execution phases by utilizing control barrier functions. Additionally, a bi-level strategy is proposed in [104] to further enhance safety and efficiency. At the upper level, the MADQN model is employed to make lane-change decisions, taking into account the collaborative nature of driving and considering the driving intentions of surrounding vehicles, enabling implicit negotiation for right-of-way. At the lower level, a right-of-way assignment model is utilized to ensure safety. A novel reward function is further proposed to encourage coordination and account for traffic impact. Moreover, Li et al. have introduced MetaDrive, a platform that focuses on safe driving by addressing generalizability, safety awareness, and multi-agent decision-making. It offers diverse driving scenarios for benchmarking single and multi-agent RL tasks [120].
@@ III-C Three-dimensional cooperation
Three-dimensional cooperation involves coordinating both longitudinal and lateral motion as well as the temporal dimension in diverse traffic scenarios, including traffic signal coordination, on-ramp merging, and unsignalized intersections, as depicted in Fig. 8. To ensure safe, efficient, and effective cooperation, CAVs need to rapidly learn longitudinal and lateral maneuvers within tight time constraints.
@@ III-C1 Traffic signal control (TSC)
Value-based MARL algorithms have been extensively investigated for TSC challenges. For instance, to alleviate the curse of dimensionality and environmental nonstationarity problems, a decentralized coordination MADQN approach is proposed in [121] to explicitly identify and dynamically adapt agent coordination needs during the learning process. Similarly, in [122], the complex TSC problem is decomposed into simpler subproblems and tackled using multiple regional agents and a centralized global agent. Each regional agent learns its RL policy for smaller regions with reduced action spaces, while the centralized global agent combines the RL contributions from various regional agents to form a final Q-function for the entire large-scale traffic grid. Notably, QMIX, an extension of VDN, was investigated in [123] for TSC, achieving promising results. Additionally, a multi-agent dueling-double-deep Q network (MAD3QN) is introduced in [124] for TSC. This framework incorporates a novel $\gamma$-reward design that includes both the original $\gamma$-reward and $\gamma$-attention-reward. By employing a spatial differentiation method for agent coordination, their approach enables decentralized control and decoupling of road networks, enhancing scalability and convergence. In [125], a cooperative group-based multi-agent Q-learning for ATSC (CGB-MATSC) is proposed. To enhance the learning ability, k-nearest-neighbor-based state representation, pheromone-based regional green-wave control mode, and spatial discounted reward design are introduced.
On the other hand, policy-based MARL algorithms have also been widely studied for TSC. In [32], a fully scalable and decentralized Multi-Agent Actor-Critic (MA2C) algorithm is introduced, which incorporates an advanced discount factor to mitigate the effects of remote agents. This approach demonstrates significant advantages over random and independent controllers. Wu et al. introduced a novel multi-agent recurrent deep deterministic policy gradient (MARDDPG) algorithm for TSC in vehicular networks [126]. They adopt a strategy of centralized training and decentralized execution, while also implementing parameter sharing among the actor networks to accelerate the training process. Additionally, LSTM networks are incorporated to leverage historical information for improved control. In [127], a MA2C approach with a feudal hierarchy concept is presented. In this approach, the traffic network is divided into several regions, each monitored by a manager agent along with traffic signal agents as workers. Managers handle high-level coordination and goal-setting for workers who control traffic signals to achieve these goals. This hierarchical structure allows for global coordination while preserving scalability.
@@
Graph-based MARL offers advantages in flexible representation, scalability, efficient message passing, a global view, interpretable representations, and adaptability, making it suitable for various multi-agent scenarios, including TSC problems [130]. In [131], the authors present CoLight, a model that employs graph-attentional MARL networks to enable communication among traffic signals, integrating temporal and spatial influences from neighboring intersections to a target intersection while establishing index-free modeling of these neighboring intersections. In [132], an Inductive Heterogeneous Graph Multi-agent Actor-Critic (IHG-MA) algorithm for multi-intersection traffic signal control is proposed, featuring two key aspects. Firstly, it utilizes an inductive heterogeneous graph (IHG) neural network for representation learning, effectively encoding both heterogeneous features and structural information of nodes and graphs to generate embeddings for previously unseen nodes and new traffic networks. Secondly, IHG-MA adopts the MA2C framework for policy learning, utilizing these embeddings to compute Q-values and policies.
To account for the spatial influence of multi-intersection traffic lights and the temporal dependency on historical traffic conditions, Wang et al. introduce a novel spatio-temporal MARL framework (STMARL) [133]. This framework captures spatio-temporal dependencies among multiple traffic lights, coordinating their control through a traffic light adjacency graph constructed based on spatial structure and integrating historical traffic records with current traffic status using a recurrent neural network. Additionally, a graph neural network model is designed to represent relationships among multiple traffic lights based on temporally-dependent traffic information. Considering the importance of shared information, relying on heuristic methods. Jiang et al. propose a new communication form called UniComm, which incorporates numerous observations gathered at one intersection and predicts their impact on neighboring intersections, enhancing communication efficiency [134]. Antonio et al. leverage the multi-agent TD3 algorithm by incorporating an LSTM to address the issue of varying observation shapes based on the number of vehicles, and then employ a training method called curriculum through self-play to enable collaborative control of CAVs at intersections [135].
@@ III-C2 On-ramps merging
Highway on-ramp merging is a challenging task for AVs, as it involves the efficient merging of on-ramp CAVs onto the through lane without collisions (see Fig. 8 (b)). In this process, CAVs on the through lane need to proactively adjust their speed, either decelerating or accelerating, to create sufficient space for on-ramp CAVs to merge safely. Simultaneously, on-ramp CAVs adapt their speed and execute lane changes promptly when it is safe to do so, in order to prevent deadlock situations [136, 35].
@@
Some studies also examine the influence of human drivers in on-ramp merging scenarios. For instance, Hu et al. introduce interaction-aware decision-making with an adaptive strategies (IDAS) approach, which utilizes a modified MA2C method and curriculum learning to train a single policy that enables an AV to navigate the road while taking into account the cooperativeness of other drivers [140]. In [35], the authors present MA2C approach that enables CAVs on both the merge lane and through the lane to collaboratively learn a policy, maximizing traffic throughput while adapting to the behavior of HDVs. Their efficient and scalable MARL framework accommodates dynamic traffic scenarios with varying communication topologies, utilizing parameter sharing, local rewards, and action masking to promote inter-agent cooperation and scalability. Additionally, a priority-based safety supervisor is introduced to reduce collision rates and expedite the training process. Similarly, a game-theoretic multi-agent planning approach called GAMEPLAN is introduced in [141], which incorporates both human drivers and CAVs in merging scenarios. This method leverages game theory and an auction mechanism to compute optimal actions for each agent, considering their driving style inferred from commonly available sensors. The algorithm assigns higher priority to more aggressive or impatient drivers and lower priority to more conservative or patient drivers, ensuring game-theoretic optimality while preventing collisions and deadlocks.
@@ III-C3 Unsignalized intersections
Intersections serve as pivotal nodes and potential bottlenecks within urban road networks. Enhancing traffic efficiency at these junctions is essential in ameliorating overall traffic throughput and alleviating congestion. Nevertheless, the management of vehicular interactions at unsignalized intersections, as illustrated in Fig. 8 (b), presents a notably intricate challenge [143]. Many conventional techniques, such as rule-based [146], planning-based [147], and single-agent RL methods [148], tend to approach the problem of intersection management as if it were a single-agent challenge, which can not capture the complex interactions among CAVs [149]. Recently, MARL has emerged as a promising and effective tool for managing traffic at unsignalized intersections.
@@
Yan et al. utilize multi-agent policy decomposition, enabling decentralized control based on local observations for any number of controlled CAVs. Remarkably, even without reward shaping, the DQN-based approach learns to coordinate CAVs to exhibit traffic signal-like behaviors, achieving near-optimal throughput with 33%-50% controlled vehicles. Furthermore, through multi-task learning and transfer learning, they demonstrate the generalizability of this behavior across different inflow rates and traffic network sizes [151].
@@ III-D Simulation platforms
MACAD (Multi-Agent Connected and Autonomous Driving), as another high-fidelity simulation platform for CAVs, adopts "partially observable Markov games" to formulate the connected autonomous driving problems with realistic assumptions and modeling. The platform provides an extensible set of CAD simulation environments that enable the research and development of Deep RL-based integrated sensing, perception, planning, and control algorithms for CAD systems with unlimited operational design domains under realistic, multi-agent settings [160]. The MACAD also provides a taxonomy of multi-agent learning environments based on the nature of tasks, the nature of agents, and the nature of the environment, which can be tuned by the researchers.
@@
Nocturne [163]
SUMO [156]
CityFlow [157]
SMARTS [34]
MetaDrive [158]
CARLA [159]
MACAD [160]
IV Discussions & Future Works

### v3

@@ III-B Two-dimensional cooperation
Compared to one-dimensional cooperation, two-dimensional cooperation addresses problems related to both longitudinal and lateral control, with its most typical application being cooperative lane changing. As shown in Fig. 7, CAVs communicate and coordinate through V2V and V2I communication channels, enabling them to make decisions and acquire effective lane-changing strategies within shared driving environments [116].
Generally, the lane-changing mathematical problem [117] can be expressed as follows:
$\min_{\mathbf{u}_{i}(t)}J$ $=\int_{t_{0}}^{t_{f}}\Bigg{[}\alpha_{1}\Big{(}\left(x_{i-1}(t)-x_{i}(t)-L_{i}-\Delta x_{\text{safe}}\right)^{2}+$ (9)
$\quad\left(x_{i}(t)-x_{i+1}(t)-L_{i+1}-\Delta x_{\text{safe}}\right)^{2}\Big{)}+\alpha_{2}\left(t_{f}-t_{0}\right)$
$\quad+\alpha_{3}\Big{(}u_{x,i}(t)^{2}+u_{y,i}(t)^{2}+\left(\frac{d}{dt}u_{x,i}(t)\right)^{2}+$
$\left(\frac{d}{dt}u_{y,i}(t)\right)^{2}\Big{)}\Bigg{]}dt$
Subject to the vehicle dynamics:
$\dot{\mathbf{x}}_{i}(t)=\mathbf{A}\mathbf{x}_{i}(t)+\mathbf{B}\mathbf{u}_{i}(t)$ (10)
where $\alpha_{1}$, $\alpha_{2}$, and $\alpha_{3}$ are the corresponding weights for safety, efficiency, and comfort, respectively. Other objectives may also be considered; $\mathbf{A}=\left[\begin{array}[]{llll}0&0&1&0\\ 0&0&0&1\\ 0&0&0&0\\ 0&0&0&0\end{array}\right],\quad\mathbf{B}=\left[\begin{array}[]{ll}0&0\\ 0&0\\ 1&0\\ 0&1\end{array}\right]$, $\mathbf{x}_{i}(t)=\left[x_{i}(t),y_{i}(t),v_{x,i}(t),v_{y,i}(t)\right]^{T}$ is the state vector for vehicle $i$; $x_{i}(t)$ and $y_{i}(t)$ are the longitudinal and lateral position of the $ith$ vehicle at time $t$; $v_{x,i}(t)$ and $v_{y,i}(t)$ are longitudinal and lateral velocity of the $ith$ vehicle at time $t$. $\mathbf{u}_{i}(t)=\left[u_{x,i}(t),u_{y,i}(t)\right]^{T}$ is the control input vector for vehicle $i$, $u_{x,i}(t)$ and $u_{y,i}(t)$ are the control input for longitudinal and lateral acceleration respectively; $L_{i}$ is the length of the $ith$ vehicle; $\Delta x_{\text{safe}}$ is the safety distance to avoid collision; $w_{\text{lane}}$ is the width of the lane.
And the constraints are expressed as:
$x_{i-1}(t)-x_{i}(t)\geq L_{i}+\Delta x_{\text{safe}}$ (11)
$x_{i}(t)-x_{i+1}(t)\geq L_{i+1}+\Delta x_{\text{safe}}$
[…]
$0\leq y_{i}(t)\leq w_{\text{lane}}$
@@
To achieve a safe, efficient, and comfortable lane-change maneuver, Hou et al. devise a decentralized cooperative lane-changing controller employing a Multi-Agent Proximal Policy Optimization (MAPPO) approach, which empowers each vehicle to independently learn and assess its policy and action rewards based on local information, while still having access to global state data. The trained policies can be effectively transferred and applied across various traffic conditions, enhancing traffic throughput from uncongested to highly congested scenarios [100]. In [103], the authors put forth a twin-vehicle cooperative driving approach, leveraging the QMIX algorithm. Their method applies RL to dynamically adapt to changing conditions on highways, aiming to find an optimal balance between autonomous decision-making and cooperative interaction among the vehicles. This algorithm allows each vehicle in the pair to independently perform lane changes and overtaking maneuvers, even in dense traffic scenarios, while maintaining a predetermined formation between them. An enhanced QMIX algorithm [118] is further proposed to enhance the flexibility and effectiveness of the collaborative lane-changing system. They implement a stable estimation method to mitigate the problem of overestimated joint Q-values among agents. This approach strikes a fine balance between maintaining formation and allowing for smooth overtaking, thus facilitating intelligent adaptation to a variety of scenarios, including heavy traffic, light traffic, and emergencies. In an effort to enhance mobility within complex traffic environments, in [119], instead of relying on relative distance and semantic maps, the authors suggest the utilization of traffic states that encapsulate the spatio-temporal interactions between neighboring vehicles. Within the MADRL framework, three prediction models, namely the transformer-based (TS-Transformer), generative adversarial network-based (TS-GAN), and conditional variational autoencoder-based (TS-CVAE) models, are developed and compared for traffic state prediction.
@@
Graph neural networks (GNNs) [120] have gained significant attention in MARL settings, particularly for cooperative lane-changing applications. Chen et al. introduce a novel algorithm based on MADQN, which combines a GCN with a deep Q-network [121]. This approach facilitates effective information fusion and decision-making within the MARL framework, thereby enhancing both safety and mobility in cooperative lane-changing scenarios. Moreover, in [122], the authors leverage GNN in conjunction with MADDPG in their work for multi-agent training, addressing the complexity of real-world inputs and aiding in congestion mitigation efforts.
@@
Safety is of paramount importance in the context of CAV scenarios. In [123], the authors have introduced a safety-enhancing actor-critic algorithm that incorporates two innovative techniques. The first technique is the ‘truncated Q-function’, which effectively leverages shared information from neighboring CAVs, ensuring scalability for large-scale CAV systems. The second technique, “safe action mapping”, provides safety guarantees throughout both the training and execution phases by utilizing control barrier functions. Additionally, a bi-level strategy is proposed in [104] to further enhance safety and efficiency. At the upper level, the MADQN model is utilized for making decisions about lane changes. This approach acknowledges the cooperative aspect of driving by factoring in the intentions of surrounding vehicles, enabling implicit negotiation for right-of-way. At the lower level, a right-of-way assignment model is utilized to ensure safety. A novel reward function is further proposed to encourage coordination and account for traffic impact. Moreover, Li et al. have introduced MetaDrive, a platform that focuses on safe driving by addressing generalizability, safety awareness, and multi-agent decision-making. It offers diverse driving scenarios for benchmarking single and multi-agent RL tasks [124].
@@ III-C Three-dimensional cooperation
Three-dimensional cooperation involves the intricate coordination of longitudinal and lateral movements along with timing, across a variety of traffic scenarios such as traffic signal coordination, on-ramp merging, and navigating through unsignalized intersections, as depicted in Fig. 8. To ensure safe, efficient, and effective cooperation, CAVs need to rapidly learn longitudinal and lateral maneuvers within tight time constraints. The three-dimensional cooperation problem can also be described by Eq. 9 - Eq. 11, with the added time constraints $t_{i,\text{ start }}\leq t\leq t_{i,\text{ end }}$.
@@ III-C1 Traffic signal control (TSC)
Value-based MARL algorithms have been extensively investigated for TSC challenges. For instance, to alleviate the curse of dimensionality and environmental nonstationarity problems, a decentralized coordination MADQN approach is proposed in [125] to explicitly identify and dynamically adapt agent coordination needs during the learning process. Similarly, in [126], the complex TSC problem is decomposed into simpler subproblems and tackled using multiple regional agents and a centralized global agent. Each regional agent learns its RL policy for smaller regions with reduced action spaces, while the centralized global agent combines the RL contributions from various regional agents to form a final Q-function for the entire large-scale traffic grid. Notably, QMIX, an extension of VDN, was investigated in [127] for TSC, achieving promising results. Additionally, Liu et al. introduce a multi-agent dueling-double-deep Q network (MAD3QN) for TSC [128]. This framework features an innovative $\gamma$-reward design, which combines the traditional $\gamma$-reward with a novel $\gamma$-attention-reward. By employing a spatial differentiation method for agent coordination, their approach enables decentralized control and decoupling of road networks, enhancing scalability and convergence. In [129], a cooperative group-based multi-agent Q-learning for ATSC (CGB-MATSC) is proposed. This method enhances the learning process by incorporating a k-nearest-neighbor approach for state representation, a pheromone-based strategy for creating regional green-wave traffic flows, and a spatially discounted reward system.
On the other hand, policy-based MARL algorithms have also been widely studied for TSC. In [34], a fully scalable and decentralized Multi-Agent Actor-Critic (MA2C) algorithm is introduced, which incorporates an advanced discount factor to mitigate the effects of remote agents. This approach demonstrates significant advantages over random and independent controllers. Wu et al. introduced the multi-agent recurrent deep deterministic policy gradient (MARDDPG) algorithm, specifically designed for TSC in vehicular networks [130]. They utilize a strategy of CTDE, enhancing the efficiency of the training process through parameter sharing among the actor networks. Furthermore, the integration of LSTM networks enables the algorithm to utilize historical data for more effective control decisions. In [131], a MA2C approach with a feudal hierarchy concept for TSC is presented. This approach segments the traffic network into several regions, each monitored by a manager agent, with traffic signal agents acting as workers. The managers are responsible for high-level coordination and setting objectives for the workers, who then adjust traffic signals to meet these objectives. This hierarchical structure fosters global coordination while ensuring the system scalability.
@@
Graph-based MARL offers advantages in flexible representation, scalability, efficient message passing, a global view, interpretable representations, and adaptability, making it suitable for various multi-agent scenarios, including TSC problems [134]. In [135], the authors present CoLight, a model leveraging graph-attentional MARL networks. This model facilitates communication among traffic signals, incorporating both temporal and spatial data from neighboring intersections to a target intersection. Crucially, it achieves index-free modeling of these neighboring intersections, enhancing its adaptability and efficiency. In [136], Yang et al. propose an Inductive Heterogeneous Graph Multi-agent Actor-Critic (IHG-MA) algorithm for controlling traffic signals across multiple intersections. This algorithm is distinguished by two main features: Firstly, it utilizes an inductive heterogeneous graph (IHG) neural network for representation learning, effectively encoding both heterogeneous features and structural information of nodes and graphs to generate embeddings for previously unseen nodes and new traffic networks. Secondly, IHG-MA adopts the MA2C framework for policy learning, utilizing these embeddings to compute Q-values and policies.
To account for the spatial influence of multi-intersection traffic lights and the temporal dependency on historical traffic conditions, Wang et al. introduce a novel spatio-temporal MARL framework (STMARL) [137]. This framework captures spatio-temporal dependencies among multiple traffic lights, coordinating their control via a traffic light adjacency graph. This graph is meticulously constructed to reflect the spatial structure of intersections and merges historical traffic data with the prevailing traffic conditions through the use of a recurrent neural network. Additionally, a GNN model is designed to represent the relationships among various traffic lights, drawing on traffic information that varies over time. Considering the importance of shared information, relying on heuristic methods. Jiang et al. propose a new communication form called UniComm, which incorporates numerous observations gathered at one intersection and predicts their impact on neighboring intersections, enhancing communication efficiency [138]. Antonio et al. leverage the multi-agent TD3 algorithm by incorporating an LSTM to address the issue of varying observation shapes based on the number of vehicles, and then employ a training method called curriculum through self-play to enable collaborative control of CAVs at intersections [139].
@@ III-C2 On-ramps merging
Highway on-ramp merging task requires the seamless integration of on-ramp CAVs into the main traffic flow without causing collisions (see Fig. 8 (b)). This complex maneuver necessitates that CAVs traveling in the through lane proactively modify their velocities—slowing down or speeding up as necessary—to create sufficient room for CAVs on the on-ramp to merge safely. Concurrently, CAVs on the ramp must regulate their velocities and make lane changes promptly in a timely manner when conditions are safe, thereby avoiding potential deadlock situations. This dual adjustment ensures a smooth and efficient merging process, critical for maintaining traffic flow and safety on highways [140, 32].
@@
Some studies also examine the influence of human drivers in on-ramp merging scenarios. For instance, Hu et al. introduce interaction-aware decision-making with an adaptive strategies (IDAS) approach, which utilizes a modified MA2C method and curriculum learning to train a single policy that enables an AV to navigate the road while taking into account the cooperativeness of other drivers [144]. In [32], the authors introduce the MA2C approach, which allows CAVs on both the merge lane and through lane to cooperatively develop a policy. This approach focuses on optimizing traffic flow while also adjusting to the dynamics of HDVs. Their efficient and scalable MARL framework accommodates dynamic traffic scenarios with varying communication topologies, utilizing parameter sharing, local rewards, and action masking to promote inter-agent cooperation and scalability. Furthermore, Chen et al. introduce a priority-based safety supervisor designed to minimize collision rates and accelerate the training process. Similarly, a game-theoretic multi-agent planning approach called GAMEPLAN is introduced in [28], which incorporates both human drivers and CAVs in merging scenarios. This method leverages game theory and an auction mechanism to compute optimal actions for each agent, considering their driving style inferred from commonly available sensors. The algorithm assigns higher priority to more aggressive or impatient drivers and lower priority to more conservative or patient drivers, ensuring game-theoretic optimality while preventing collisions and deadlocks.
@@ III-C3 Unsignalized intersections
Intersections serve as pivotal nodes and potential bottlenecks within urban road networks. Enhancing traffic efficiency at these junctions is essential in ameliorating overall traffic throughput and alleviating congestion. Nevertheless, the management of vehicular interactions at unsignalized intersections, as illustrated in Fig. 8 (b), presents a notably intricate challenge [146]. Many conventional techniques, such as rule-based [149], planning-based [150], and single-agent RL methods [151], tend to approach the problem of intersection management as if it were a single-agent challenge, which can not capture the complex interactions among CAVs. Recently, MARL has emerged as a promising and effective tool for managing traffic at unsignalized intersections.
@@
Yan et al. utilize a multi-agent policy decomposition strategy, which allows for decentralized control using local observations across an arbitrary number of CAVs. Notably, without the need for reward shaping, the DQN-based method learns to coordinate CAVs in a way that mimics traffic signal behaviors, attaining nearly optimal throughput with control over 33% to 50% of the vehicles. Furthermore, through multi-task learning and transfer learning, they demonstrate the generalizability of this behavior across different inflow rates and traffic network sizes [153].
@@ III-D Simulation platforms
MACAD (Multi-Agent Connected and Autonomous Driving)platform serves as a high-fidelity simulation environment for CAVs, utilizing "partially observable Markov games" to accurately depict the challenges of connected autonomous driving with realistic assumptions and models. This platform offers a comprehensive range of CAD simulation settings, facilitating the exploration and advancement of Deep RL strategies in integrated sensing, perception, planning, and control for CAD systems across diverse operational design domains, under realistic multi-agent conditions, as highlighted by [163]. Furthermore, MACAD categorizes multi-agent learning scenarios according to task types, agent characteristics, and environmental factors, providing researchers with customizable options.
@@
Nocturne [166] - - - -
SUMO [159] - - -
CityFlow [160] - - - -
SMARTS [35] - - - -
MetaDrive [161] - -
CARLA [162] -
MACAD [163] -
IV Discussion & Future Work